What molar ratio of sodium cyanide to hydrocyanic acid is necessary to prepare a buffer with pH = 9.72? Ka of HCN = 4.9 x 10-10
Answer #1

use the hander son equation

pH = pKa + log(CN-/HCN)

first convert Ka to Pka

pKa = -log(Ka) = -log(4.9 x 10-10) = 9.31

pH is given as 9.72

now put all these values in the above equation

9.72 = 9.31 +  log(CN-/HCN)

log(CN-/HCN) = 9.72 - 9.31

log(CN-/HCN) = 0.41

remove the log means

[CN-/HCN] = 100.41 = 2.57